Wilmington Company has two manufacturing departments—Assembly and Fabrication. It considers all of its manufacturing overhead costs to be fixed costs. The first set of data that is shown below is based on estimates from the beginning of the year. The second set of data relates to one particular job completed during the year—Job Bravo.

Estimated Data Assembly Fabrication Total
Manufacturing overhead costs $ 7,800,000 $ 8,400,000 $ 16,200,000
Direct labor-hours 150,000 90,000 240,000
Machine-hours 60,000 300,000 360,000
Job Bravo Assembly Fabrication Total
Direct labor-hours 31 23 54
Machine-hours 23 26 49

Required:

1. If Wilmington used a plantwide predetermined overhead rate based on direct labor-hours, how much manufacturing overhead would be applied to Job Bravo?

2. If Wilmington uses departmental predetermined overhead rates with direct labor-hours as the allocation base in Assembly and machine-hours as the allocation base in Fabrication, how much manufacturing overhead would be applied to Job Bravo?

(Round your intermediate calculations to 2 decimal places.)
Answer #1

1) Overhead rate on direct labor hours = 162000000/240000 = 67.50 per hour

Overhead applied to Job Bravo = 54*67.50 = 3645

2) Assembly overhead rate = 7800000/150000 = 52 per DLH

Fabrication overhead rate = 8400000/300000 = 28 per MH

Overhead applied to Job Bravo = 31*52+28*26 = 2340